OpenStudy (anonymous):

The table shows the height of a tree as it grows. What equation in slope-intercept form gives the tree's height at any time?

OpenStudy (anonymous):

The table shows the height of a tree as it grows. What equation in slope-intercept form gives the tree's height at any time? Time Height (months) (inches) 2 14 4 23 6 32 8 41 y = 9x + 2 y = 9/2x + 5 y = 9/2x y = 5x + 9/2

OpenStudy (kropot72):

The slope m of the graph of height against time is found from the formula: \[m=\frac{y _{1}-y _{2}}{x _{1}-x _{2}}=\frac{14-41}{2-8}=you\ can\ calculate?\]

OpenStudy (anonymous):

I'm still kind of confused I'm sorry :(

OpenStudy (kropot72):

We are looking for a straight line equation of the form y = mx + b where m is the slope and b is the intercept on the y axis. The first step is to use the given data to find the slope. So as I asked before can you calculate the value of the following: \[m=\frac{14-41}{2-8}=?\]